# Extreme events in locally coupled bursting neurons

### v1

@@ Abstract
We report a new mechanism through which extreme events with a dragon king-like distribution emerge in a network of locally coupled Hindmarsh-Rose bursting neurons. We establish and substantiate the fact that depending on the choice of initial conditions, the neurons in the network are divided into clusters and whenever these clusters are phase synchronized intermittently, extreme events originate in the collective observable. This mechanism, which we name as intermittent cluster synchronization is proposed as the new precursor for the generation of extreme events in this system. These results are also true for electrical diffusive coupling. The distribution of the local maxima shows long tailed non-Gaussian while the interevent interval follows the Weibull distribution. The goodness of fit are corroborated using probability-probability plot and quantile-quantile plot. These extreme events become rarer and rarer with the increase in the number of different initial conditions.
@@ I Introduction
Despite the lack of an universally accepted definition, EE are classified by their abnormal characteristics. Statistical indicators like the significant height $H_{s}$ help to distinguish EE from other regular events [5, 6, 7]. In a dynamic sense, EE are described as a sudden shift of a variable from a normal state to a high-amplitude state. A long-tail non-Gaussian probability distribution is typically seen when analyzing the amplitudes of such dynamic observables, indicating the presence of infrequent large-scale events [8]. Sometimes dragon-king distribution has also been observed in the distribution patterns [1]. In these cases, most small to medium-scale events (city populations, earthquake sizes, stock price fluctuations, etc.) follow a power law, except for sporadic huge events that deviate as outliers. On the other hand, it has been shown that the significant mechanisms behind these abrupt transitions in isolated systems are interior crisis [9, 10, 11, 12, 13, 14], Pomeau–Manneville (PM) intermittency [9, 15], noise-induced intermittency [16], sliding bifurcation [17], alternation in size of the chaotic attractor [12] and strange nonchaotic attractor [18, 19]. Sudden expansions of chaotic attractors through interior crises lead to EE in complex systems [20, 21, 22, 23] as well. Current research also focus on exploring other potential causes of EE both in isolated and complex systems [8, 17, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 18, 19, 34, 35, 36, 37]. Especially, in complex systems, a network of Fitzhugh-Nagumo neurons with global and small world coupling topology produces EE [20, 21]. Similarly, EE also evolve in networks comprising of Josephson junctions [29], Liénard systems [29], Stuart-Landau oscillators [38], coupled chaotic maps [39], moving agents [40], neuronal maps that are diffusively coupled locally and nonlocally [41]. These rare events manifest in complex coupled systems as a result of the excitability of proto events [20], instability of out-of-phase synchronization [27], the interplay of attractive and repulsive coupling [40], the interplay between degree distribution and repulsive coupling [42], and counter-rotating chaotic attractors [38]. These are the very few mechanisms that have been determined so far.
@@
In this article, we carry out the dynamical analysis of EE and determine its precursor in a locally coupled network of HR neurons whose individual nodes are nonlinearly (through chemical synapse) connected with each other. From our analysis, we unveil an important and a new mechanism that intermittent cluster synchronization among the neurons in the network produces EE. For the chosen number of the different initial conditions (ICs), correspondingly neurons in the network spilt into many clusters. These clusters evolve desynchronously most of the time and occasionally synchronize for a short duration. During such an intermittent synchronization, EE are observed in the collective observable of the network. The rarity of the EE are found to increase with increase in the number of ICs. Typically, intermittent synchrony is observed in brain neuron networks and the EE emerging in them are deeply related to epilepsy (like epileptic seizures) [53] when a highly detrimental disorder happening in the human brain collapsing the individual. Present study and the mechanism found out of this work will serve as an important milestone in the understanding of dynamical mechanisms behind epileptic seizures.
We present our work as follows. In Sec. II, we describe the model that we take for investigation, discuss the emergence of EE in particular network topology, detail the statistics followed by them and discuss about the spiking to bursting transition. In Sec. III, we demonstrate the mechanism behind the events. The Robustness of EE to the ICs is studied in Sec. IV. Finally in Sec. V, we summarize and conclude.
@@ II Locally Coupled Network
The Hindmarsh-Rose model [43] is well known for the chaotic and different types of bursting behaviour [46]. The mathematical form is expressed as
@@
where the variable $x$ represents the membrane potential, $y$ and $z$ represent the transport of ions in the fast and slow channel respectively. The variable $y$ controls the variation of the fast current, and the variable $z$ represents the variation of the slow current. The speed is controlled by the small parameter $c$. The parameter $I$ gives the external current that enters the neuron. $x_{r}$ is the parameter that controls the delaying and advancing activation of the slow current in the neuron.
Figure 1: Schematic diagram of the $N$ unidirectionally coupled systems. Each numbering corresponds to a node each evolving according to Eq. (1).
Now let us consider $N$ identical copies of HR neurons which are coupled unidirectionally as shown in the schematic Fig. 1. Modifying Eq. (1) for this schematic, we get
@@
where $i$ is the number of neurons in the system with the boundary condition ($x_{N+1}=x_{1}$) and $N$ is the total number of neurons. In our analysis, we take $N=100$. With this boundary condition, the chain of locally coupled oscillators looks like a ring, as visible from Fig. 1. $i^{th}$ oscillator is coupled to the $(i+1)^{th}$ oscillator nonlinearly through the sigmoidal function ($\Gamma(x)$): a type of coupling function which is commonly used to represent the chemical synaptic interaction [47]. The mathematical form of the sigmoidal function is given by [46]
@@
The parameters $\lambda$ and $\theta$ in Eq. (3) correspond to the slope of the function and the synaptic threshold respectively. Throughout our study, we fix the values of the system parameters at $a=1,b=3,c=1,d=5,x_{r}=-1.6,r=0.01,s=5,v_{s}=2$, $\lambda=10$, $\theta=-0.25$. $I$ is the current given to the neurons and $k$ the coupling strength between the neurons is taken as the bifurcation parameter.
Here, we investigate two cases, namely inhibitory where $k$ is negative and excitatory where $k$ is positive. Subsequently, in the following, we will discuss EE resembling dragon kings appearing in the system (2). Equation (2) is integrated numerically using the fifth-order Runge-Kutta-Fehlberg integration scheme with a fixed step length of $\Delta t=0.01$. For our numerical analysis, a set of initial conditions is used to initialize the state variables. Among the $N=100$ neurons, odd neurons are set to the initial conditions (IC’s): $x_{odd}=1,y_{odd}=0.5,z_{odd}=1$, and the even neurons are set to the IC’s: $x_{even}=0.5,y_{even}=1,z_{even}=0.5$. In our study, we consider the mean of the $x$ variable $X_{m}=\sum_{i=1}^{N}x_{i}$ as the observable. Similarly $Y_{m}=\sum_{i=1}^{N}y_{i}$ and $Z_{m}=\sum_{i=1}^{N}z_{i}$.
@@ II.1 Emergence of extreme events
We start our analysis by using the peak over threshold method [8] to identify EE appearing in the system. In this connection, we define the events that are larger than a predefined threshold level ($H_{s}$) as EE [8]. The threshold level is estimated using
@@
where $\mu$ is the mean value and $\sigma$ is the standard deviation of all the peaks of the collective observable $X_{m}$. In our case, we chose $n=6$.
We perform numerical integration of system (2) up to $5\times 10^{7}$ iterations. Leaving out $2\times 10^{6}$ iterations as transients and then we collect and plot in Fig. 2(a), the probability of the EE ($P_{EE}$) in the $(k-I)$ parameter space. The calculation is carried out using the relation
$P_{EE}=\frac{\text{Number of peaks crossing the threshold}}{\text{Total number of peaks}}.$ (5)
@@
The colour bar in Fig. 2a displays the $P_{EE}$, where lime green (grey) represents the highest probability and dark blue (black) represents the lowest probability. In the white regions, no EE occur. Further observation of Fig. 2a reveals that more regions from the inhibitory side are prone to produce EE compared to the areas from the excitatory direction. Between the zones $I=3.6$ and $3.85$, we can see a short straight region in the coupling strength $k=0.04$ and $0.05$ where EE occur more than other regions. We investigate these EE’s birth and other statistical properties by discerning the region covered by the constant ordinate $I=4$. When we look at the region $I=4$ in Fig. 2(a), EE appear almost symmetrically in both directions of the coupling strength. To confirm this, we plot in Fig. 2(b), $P_{EE}$ with respect to the coupling strength $k$ at $I=4$. Increasing $k$, we can observe EE appearing in the system in both inhibitory and excitatory domains of the coupling strength $k$. Figure 2(b) is also obtained for $5\times 10^{7}$ iterations after leaving out transients of $2\times 10^{6}$ iterations. From Fig. 2(b), we can see that EE occur in the inhibitory regions from $k=-0.23$ to $k=-0.06$ and from $k=0.06$ to $k=0.17$ in the excitatory regions.
@@ II.2 Statistical showcase
In Fig. 3(a), we plot the probability distribution function of peaks by collecting all the local maximum. From the plot, we can observe a slight hump at the back. From this, we can infer that the distribution of local maximum follows a dragon king-like distribution. Hence EE occurring in the system (2) are dragon king-like EE. Further, in Fig. 3(b), we plot the Gaussian distribution of the peaks where the distribution can be found to have a very long tail. This tail represents the EE and the probability of these tail events is near to zero but still non-zero. This represents the rarity of these events. Despite being rare, the consequences of these events are catastrophic. Next, we plot the inter-event interval (IEI) in Fig. 3(c). We find that Weibull Distribution fits very well with the IEI. These fitted values were acquired using the python package “scipy.stats” where we acquired the values of the shape and scale parameters respectively as $1.00286$ and $3522.28717$. We used the Freedman-Diaconis rule to set the number of bins to our simulated data sets [39]. Figures 3(d) and 3(e) represent the probability-probability (PP) and quantile-quantile (QQ) plots which decide the quality and correlation of the fit with the actual data sets. In Fig. 3(d), blue solid line is the cumulative distribution function (CDF) of the stimulated data sets and the red dashed line is the CDF of the fit. From this, we can observe most of the times the fitted and stimulated data sets correlate with each other which shows the quality of the fit. Figure 3(e) is the QQ plot. In this, the simulated data sets are considered as the cut points which divide the range of the continuous interval with equal probabilities. We used ”PROBPLOT FUNCTION IN PYTHON” to plot the QQ plot. This function essentially estimates a probability plot for the given data with respect to the quantiles of the utilized distribution.

### v2

# Intermittent cluster synchronization in a unidirectional ring of bursting neurons

@@ Abstract
We report a new mechanism through which extreme events with a dragon king-like distribution emerge in a network of unidirectional ring of Hindmarsh-Rose bursting neurons interacting through chemical synapses. We establish and substantiate the fact that depending on the choice of initial conditions, the neurons are divided into different clusters. These clusters transit from a phase-locked state (anti-phase) to phase synchronized regime with increasing value of the coupling strength. Before attaining phase synchronization, there exists some regions of the coupling strength where these clusters are phase synchronized intermittently. During such intermittent phase synchronization, extreme events originate in the mean-field of the membrane potential. This mechanism, which we name as intermittent cluster synchronization, is proposed as the new precursor for the generation of emergent extreme events in this system. These results are also true for diffusive coupling (gap junctions). The distribution of the local maxima of the collective observable shows a long tailed non-Gaussian while the interevent interval follows the Weibull distribution. The goodness of fit are corroborated using probability-probability plot and quantile-quantile plot. This intermittent phase synchronization becomes rarer and rarer with the increase in the number of clusters of initial conditions.
@@ I Introduction
Despite the lack of an universally accepted definition, EE are classified by their abnormal characteristics. Statistical indicators like the significant height $H_{s}$ help to distinguish EE from other regular events [5, 6, 7]. In a dynamic sense, EE are described as a sudden shift of a variable from a normal state to a high-amplitude state. A long-tail non-Gaussian probability distribution is typically seen when analyzing the amplitudes of such dynamic observables, indicating the presence of infrequent large-scale events [8]. Sometimes dragon-king distribution has also been observed in the distribution patterns [1]. Extreme events also called sometimes as “Black swans” which arise as a continuous process in the tail of the power-law distribution. In contrast, “Dragon-kings” is not a result of a continuous process and hence it does not fall in the tail of a power-law distribution, rather those events deviate from the power-law distribution and follows a separate non-Gaussian distribution. Most small to medium-scale events (city populations, earthquake sizes, stock price fluctuations, etc.) follow a power law, except for sporadic huge events that deviate as outliers. On the other hand, it has been shown that the significant mechanisms behind these abrupt transitions in isolated systems are interior crisis [9, 10, 11, 12, 13, 14], Pomeau–Manneville (PM) intermittency [9, 15], noise-induced intermittency [16], sliding bifurcation [17], alternation in size of the chaotic attractor [12] and strange nonchaotic attractor [18, 19]. Sudden expansions of chaotic attractors through interior crises lead to EE in complex systems [20, 21, 22, 23] as well. Current research also focus on exploring other potential causes of EE both in isolated and complex systems [8, 17, 24, 25, 3, 26, 27, 28, 29, 30, 31, 32, 18, 19, 33, 34, 35, 36]. Especially, in complex systems, a network of Fitzhugh-Nagumo neurons with global and small world coupling topology produces EE [20, 21]. Similarly, EE also evolve in networks comprising of Josephson junctions [28], Liénard systems [9], Stuart-Landau oscillators [37], coupled chaotic maps [38], moving agents [39], neuronal maps that are diffusively coupled locally and nonlocally [40]. These rare events manifest in complex coupled systems as a result of the excitability of proto events [20], instability of out-of-phase synchronization [26], the interplay of attractive and repulsive coupling [39], riddled basing [41], on-off intermittency [38, 39], attractor bubbling [25], the interplay between degree distribution and repulsive coupling [31], and counter-rotating chaotic attractors [37]. These are the very few mechanisms that have been determined so far.
@@
In this article, we carry out the dynamical analysis of EE and determine its precursor in a locally coupled network of HR neurons whose individual nodes are nonlinearly (through chemical synapse) connected with each other. Previously, it is known that extreme events appear near synchronization manifold or are associated with intermittent synchronization/desynchronization. The relating scenarios are (i) intermittent bubbling of the transverse variable, where the transverse variable intermittently bubbles out from the invariant manifold [25], (ii) interruption of the synchronized behavior pertaining to slips in the phase difference leading to an imperfect phase synchronization [20, 21], (iii) excursion of the trajectories in the transverse direction to the synchronization manifold [22], (iv) instability of out-of-phase synchronization, where an instability arises in the desynchronization manifold causing the trajectories of two oscillators to intermittently synchronize [26], and (v) on-off intermittency manifesting as a result of intermittent desynchronization of the oscillators [38]. Along these lines, we unveil an important and a new mechanism that intermittent cluster synchronization among the neurons in the network produces EE. From our analysis, we find that for the chosen number of the clusters of initial conditions (ICs), correspondingly neurons in the network split into that many clusters. These clusters evolve desynchronously most of the time and occasionally synchronize for a short duration. During such an intermittent synchronization, EE are observed in the collective observable of the network. This scenario proposed in this article is a novel mechanism and is reported for the first time to the best of our knowledge. In general, systems with inherent symmetry is found to transit from antiphase to in-phase synchronization upon varying the coupling strength and symmetry in the initial conditions might lead to results that are not generalizable to more complex scenarios [52, 53, 54, 55, 56]. In the present case, intermittently, in time for a particular value of the coupling strength, intermittent cluster synchronization occurs leading to extreme events in the mean field variable. This occurs irrespective of whether symmetry in the clusters of ICs is present in the system or not. Further, the rarity of the EE is found to increase with increase in the number of ICs. Typically, intermittent synchrony is observed in brain neuron networks and the EE emerging in them are deeply related to epilepsy (like epileptic seizures) [57, 58, 59] when a highly detrimental disorder happening in the human brain collapsing the individual. Present study and the mechanism found out of this work will serve as an important milestone in the understanding of dynamical mechanisms behind epileptic seizures.
We present our work as follows. In Sec. II, we describe the model that we take for investigation, discuss the emergence of EE in particular network topology, detail the statistics followed by them and discuss about the spiking to bursting transition. In Sec. III, we demonstrate the mechanism behind the events. The Robustness of EE to the ICs is studied in Sec. IV. Finally in Sec. V, we summarize and conclude the findings.
@@ II Locally Coupled Network
The HR neuronal model [42] is well known for the chaotic and different types of bursting behaviour [45]. Despite being a non-physiological model (due to the cubic nonlinearity), HR model serves to be a prototypical model for bursting and helping us to qualitatively understand the nature of neuron’s dynamics in an easier way when compared to the other physiological models. The mathematical form of the HR model is expressed as
@@
where the variable $x$ represents the membrane potential, $y$ and $z$ represent the transport of ions in the fast and slow channel respectively. The variable $y$ describes the dynamics of fast current, while the variable $z$ characterizes the dynamics of the slow current. The speed of the slow variable $z$ is controlled by the parameter $r$. The parameter $I$ gives the external current that enters the neuron. $x_{r}$ is the parameter that controls the delaying and advancing activation of the slow current in the neuron. Other parameters, namely $a,~{}b,~{}c,~{}d,~{}\text{and},~{}S$ are positive coefficients.
Figure 1: Schematic diagram of the $N$ unidirectionally coupled systems. Each numbering corresponds to a node, each evolving according to Eq. (1).
Now let us consider $N$ identical copies of HR neurons [45, 60] which are coupled in a unidirectional ring topology as shown in the schematic Fig. 1. Modifying Eq. (1) for this schematic, we get
@@
where $i$ is the number of neurons in the system with the boundary condition ($x_{N+1}=x_{1}$) and $N$ is the total number of neurons. In our analysis, we take $N=100$. With this boundary condition, the chain of locally coupled oscillators looks like a ring, as visible from Fig. 1. $i^{th}$ oscillator is coupled to the $(i+1)^{th}$ oscillator nonlinearly through the sigmoidal function ($\Gamma(x)$): a type of coupling function which is commonly used to represent the chemical synaptic interaction [46]. The mathematical form of the sigmoidal function [45] is given by
@@
The parameters $\lambda$ and $\theta$ in Eq. (3) correspond to the slope of the function and the synaptic threshold respectively. Throughout our study, we fix the values of the system parameters at $a=1,b=3,c=1,d=5,x_{r}=-1.6,r=0.01,s=5,v_{s}=2$, $\lambda=10$, $\theta=-0.25$. $I$ is the current given to the neurons and $k$, the coupling strength between the neurons, is taken as the bifurcation parameter. For the chosen set of parameters, the uncoupled neuron exhibits three periodic dynamics.
Here, we investigate two cases, namely inhibitory coupling where $k$ is negative and excitatory coupling where $k$ is positive. Subsequently, in the following, we will discuss EE resembling dragon kings appearing in the system (2). Equation (2) is integrated numerically using the fifth-order Runge-Kutta-Fehlberg integration scheme with a step length $\Delta t=0.01$ [61]. We have considered the HR neuron model since this model gives a variety of dynamics by changing the parameter I. Since, the understanding of collective behaviours in complex neuronal networks is a challenging task, so for simplicity, we have chosen ring type of HR neurons where the coupling is unidirectional. Since, the ring type of network topology is symmetric, we can consider two distinct ICs, odd and even. For our numerical analysis, a set of initial conditions is used to initialize the state variables. Among the $N=100$ neurons, odd neurons ($i=1,3,...,99$) are set to the initial conditions (IC’s): $x_{odd}=1,y_{odd}=0.5,z_{odd}=1$, and the even neurons ($i=2,4,...,100$) are set to the IC’s: $x_{even}=0.5,y_{even}=1,z_{even}=0.5$. We consider the mean of the $x$ variable $X_{m}=\frac{1}{N}\sum_{i=1}^{N}x_{i}$ as the observable. Similarly $Y_{m}=\frac{1}{N}\sum_{i=1}^{N}y_{i}$ and $Z_{m}=\frac{1}{N}\sum_{i=1}^{N}z_{i}$.
@@ II.1 Emergence of extreme events
We start by using the peak over threshold method [8] and classify the events that are larger than a predefined threshold level ($H_{s}$) as EE. The threshold level is estimated using
@@
where $\mu$ is the mean value and $\sigma$ is the standard deviation of the collective observable $X_{m}$. In our case, we chose $n=6$.
We perform numerical integration of system (2) up to $5\times 10^{7}$ iterations of which $2\times 10^{6}$ iterations were left out as transients and then the we plot the probability of the EE ($P_{EE}$) in the $(k-I)$ parameter space in Fig. 2(a). The calculation is carried out using the relation
$P_{EE}=\frac{\text{Number of points crossing the threshold}}{\text{Total number of points}}.$ (5)
@@
Figure 3: Bifurcation dynamics of the collective observable with respect to the coupling strength $k$ at $I=4.0$.
The colour bar in Fig. 2a displays the $P_{EE}$, where lime green (grey) represents the highest probability and dark blue (black) represents the lowest probability. In the white regions, no EE occur. Further observation of Fig. 2a reveals that more regions from the inhibitory coupling side are prone to produce EE compared to the areas from the excitatory coupling direction. Between the zones $I=3.6$ and $3.85$, we can see a short straight region in the coupling strength $k=0.04$ and $0.05$ where EE occur more than other regions. We investigate these EE’s birth and other statistical properties by discerning the region covered by the constant ordinate $I=4$. When we look at the region $I=4$ in Fig. 2(a), EE appear almost symmetrically in both directions of the coupling strength. To confirm this, we plot in Fig. 2(b), $P_{EE}$ with respect to the coupling strength $k$ at $I=4$. Increasing $k$, we can observe EE appearing in the system in both inhibitory and excitatory coupling strengths $k$. Figure 2(b) is also obtained for $5\times 10^{7}$ iterations after leaving out transients of $2\times 10^{6}$ iterations. From Fig. 2(b), we can see that EE occur for inhibitory coupling from $k=-0.23$ to $k=-0.06$ and from $k=0.06$ to $k=0.17$ for excitatory coupling.
In Fig. 3, the bifurcation plot of the collective observable $X_{m}$ with respect to $k$ at $I=4$ is presented. We can observe that the system transits from periodic to chaotic nature upon increasing the coupling strength. This is true for both inhibitory as well as excitatory couplings. Upon comparing Fig. 2(b) and Fig. 3, we can observe that while increasing the value of the coupling strength $k$ from $-0.5$, whenever EE arises in the system (i.e., non-zero $P_{EE}$ in Fig. 2(b)), the collective dynamics exhibits chaos (Fig. 3) and occasional in-phase emerges between two clusters. For the periodic regime in the bifurcation diagram in Fig. 3, the probability of EE is zero due to the antiphase synchronization between the two clusters. The similar phenomena is also true for diffusive coupling case as well, which is discussed in the Appendix. The detailed mechanism of the emergence of EE is discussed in Sec. III.
Now, to see the local dynamics of the neurons, we plot the bifurcation diagram of individual neuron (say, 1) of system (2) respectively for three different values of the coupling strength $k=-0.4$, $k=-0.17$ and $k=0.4$ in Figs. 4(a-c). We can observe the transition from periodic to the chaotic states upon increasing the coupling strength $k$. At $k=-0.4$, the dynamics is completely periodic whereas, at $k=-0.17$ and $k=0.4$, the periodic states are interspersed by chaotic states through period doubling route. Individual neurons may exhibit periodic or chaotic dynamics, but for appropriate coupling strengths, intermittent cluster synchronizations occurs among the clusters of neurons.
Figure 4: Bifurcation diagrams of the node-$1$ with respect to $I$ at the coupling strength (a) $k=-0.4$, (b) $k=-0.17$, (c) $k=0.4$.
@@ II.2 Statistical showcase
In Fig. 5(a), we plot the probability distribution function of peaks by collecting the local maxima. From the plot, we can observe a slight hump at the back. From this, we can infer that the distribution of local maxima follows a dragon king-like distribution. Hence EE occurring in the system (2) are dragon king-like EE. Further, in Fig. 5(b), we plot the non-Gaussian distribution of the peaks where the distribution can be found to have a very long tail. This tail represents the EE and the probability of these tail events is near to zero but still non-zero. This represents the rarity of these events. Despite being rare, the consequences of these events are catastrophic. Dragon-kings basically exhibit clustering property. That is, very high probability for its occurrences beyond the threshold. This property makes them to significantly deviate from the power-law or any other normal distribution. In a system like network of neurons, where EE represent epileptic seizures, such a rare and high magnitude tail event is lethal. Next, we plot the inter-event interval (IEI) in Fig. 5(c). We find that Weibull Distribution fits very well with the IEI. These fitted values were acquired using the python package “scipy.stats” where we acquired the values of the shape and scale parameters respectively as $1.00286$ and $3522.28717$. We used the Freedman-Diaconis rule to set the number of bins to our simulated data sets [38]. Figures 5(d) and 5(e) represent the probability-probability (PP) and quantile-quantile (QQ) plots which decide the quality and correlation of the fit with the actual data sets. In Fig. 5(d), blue solid line is the cumulative distribution function (CDF) of the simulated data sets and the red dashed line is the CDF of the fit. From this, we can observe most of the times the fitted and simulated data sets correlate with each other which shows the quality of the fit. Figure 5(e) is the QQ plot. In this, the simulated data sets are considered as the cut points which divide the range of the continuous interval with equal probabilities. We used ”PROBPLOT FUNCTION IN PYTHON” to plot the QQ plot. This function essentially estimates a probability plot for the given data with respect to the quantiles of the utilized distribution.